OpenStudy (anonymous):

-5/6,-4/9,-3/5,-4/7 --- List from Least to Greatest

OpenStudy (anonymous):

May you help @pooja195

OpenStudy (tkhunny):

Well, do it. Sometimes, a nice decimal equivalent helps. -5/6 = -0.833333... -4/9 = -0.444444... -3/5 = -0.600000... -4/7 = -0.571428... More clear?